A system at equilibrium is represented by the following chemical equation: N₂(g) + 3H₂(g) <--> 2NH₃(g) At a given temperature, 1.40 moles of N₂, 0.888 moles of H₂, and 3.17 * 10⁻⁸ moles of NH₃ are in the 4.00 L vessel that reaches equilibrium. Calculate the value of K for the system.
1
Expert's answer
2020-06-08T15:14:59-0400

Concentration of N2 at equilibrium =

1.40 / 4 = 0.35 mole / litre.

Concentration of H2 at equilibrium =

0.888 / 4 = 0.222 mole / litre.

Concentration of NH3 at equilibrium =

3.17 × 10 -8 / 4 = 0.7925 × 10-8 mol / litre.

= 7.9 × 10 - 9 mol/ litre.


Equilibrium CONSTANT for reaction =

(Concentration of NH3)2

/(Concentration of N2)×(Concentration of H2)3.


So, equilibrium constant =

( 7.9 × 10-9 )2 / (0.35 ) × (0.222)3 .


Equilibrium CONSTANT = 16297.7 × 10-18 ..

Equilibrium Constant = 1.6297733 × 10-14.....